OpenStudy (steve816):

Please help I am bad at this A stone rolling horizontally off a 35-meter high cliff lands 25 meters away. How long did it spend in the air and what was its initial velocity?

OpenStudy (blazeryder):

@osprey

OpenStudy (osprey):

Please help I am bad at this A stone rolling horizontally off a 35-meter high cliff lands 25 meters away. How long did it spend in the air and what was its initial velocity? At launching point, VERTICAL velocity is zero, and HORIZONTAL velocity is UNKNOWN. Got to calculate HOW LONG the stone was in the air for. Kinematic equation. Applied to the VERTICAL MOTION S=ut+1/2 at^2 35=1/2 10 t^2 root(70/10)=time of flight about 2.6 seconds. a is acceleration, which is gravity, which is about 10. s is distance. u is initial (vertical, here) speed. guess what t is. The vertical motion and the horizontal motion are INDEPENDENT of each other, but they do COMBINE to make a rather lovely trajectory (if you're a golfer) or parabola if it's a physics question. HOPE the above is right and helps. It should be a start. http://perendis.webs.com

OpenStudy (steve816):

Thank you so much @osprey

OpenStudy (vanesaretana):

Yes, I agree @osprey

OpenStudy (osprey):

the stone was in the air for 2.6 seconds. In that time it covered a HORIZONTAL distance of 25 metres. The speed IT WOULD NEED TO HAVE would be speed = distance/time =25/2.6 = 9.45 m/s